the distance travelled by an object is given by x= at +bt^2/ c+a where t is time ad , b, c are constants find the dimensions of b, c
 

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
x = at + bt^2/c+a 
[L] = a[T] + b[T]^2/c+ a 
Measurement of b is [LT-2] 
Measurement of c is [LT2].
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Zubeen Naqvi
one year ago
X=at+bt²/(c+a) 
Dimension of at =dimension of x 
a[T]=[L]
a=[LT^-1]
Dimension of bt²/(c+a)=x
b[T²]/(c+a)=[L]
b/(c+a)=[LT^-2]
To add c and a their dimension should be same 
C=[LT^-1] 
b/[LT^-1]=[LT^-2]
b=[L²T^-3]